OpenStudy (anonymous):

another problem :( solve the equation: (x-3)(x+6)=10

OpenStudy (anonymous):

I'm not sure if this is right, so don't get mad if it isn't... but this is what i got: (9x-3)(x+6)=10 -6 -6 (x-3)x=4 x^x-3=4 +3+3 x^x=7 x=sq rt. of 7 x~2.6

OpenStudy (anonymous):

this is wrong..

OpenStudy (anonymous):

I thought so

OpenStudy (anonymous):

(x-3)(x+6)=10 => x^2+3x-28=0 =>(x+7)(x-4)=0 so, x=4, -7

OpenStudy (anonymous):

thanks :D my math homework is all onine, so it tells me if im right or wrong right away, and that was the correct answer :D thanks :D

OpenStudy (anonymous):

welcome..